### v2

# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

Abstract
Photonics represents one of the most promising approaches to large-scale quantum computation with millions of qubits and billions of gates, owing to the potential for room-temperature operation, high clock speeds, miniaturization of photonic circuits, and repeatable fabrication processes in commercial photonic foundries. We present an end-to-end architecture for fault-tolerant continuous-variable (CV) quantum computation using only passive on-chip components that can produce photonic qubits above the fault tolerance threshold with probabilities above 90%, and encodes logical qubits using physical qubits sampled from a distribution around the fault tolerance threshold. By requiring only low photon number resolution, the architecture enables the use of high-bandwidth photodetectors in CV quantum computing. Simulations of our qubit generation and logical encoding processes show a Gaussian cluster squeezing threshold of 12 dB to 13 dB. Additionally, we present a novel magic state generation protocol which requires only 13 dB of cluster squeezing to produce magic states with an order of magnitude higher probability than existing approaches, opening up the path to universal fault-tolerant quantum computation at less than 13 dB of cluster squeezing [^1] Approved for Public Release: Distribution Unlimited.
I Introduction
Quantum computation promises the ability to exponentially speed up certain problems intractable to classical computers including simulations of large quantum systems, an observation first formalized by Richard Feynman Feynman (1982). Noisy intermediate scale quantum computers Preskill (2018) with $\approx 100$ physical qubits and $\approx 100$ to $1000$ quantum gates have been demonstrated on a variety of physical platforms Debnath et al. (2016); Kandala et al. (2017); Henriet et al. (2020); O’Malley et al. (2016), but it is widely expected that the computers must scale to millions of physical qubits and billions of operations to be useful for commercially valuable quantum simulations Kim et al. (2022); Gidney and Ekerå (2021); Litinski (2023). Matter based approaches face hurdles to get to such a computational volume, owing to the hardware requirements of low temperatures, slow gate rates, or lossy conversions between matter and photons Meesala et al. (2024); Kandala et al. (2017); O’Malley et al. (2016); Henriet et al. (2020); Saffman (2016). Photonics provides a promising path towards large-scale quantum computing with scalable generation and manipulation of qubits at room temperature, high-speed measurements, miniaturization using photonic integrated circuits, and repeatable low-loss mass manufacturing at commercial foundries.
Photonic quantum information can be encoded into either discrete quantum degrees of freedom of individual single photons, known as the discrete variable (DV) approach Kok et al. (2007); Bartolucci et al. (2021), or in continuous-valued orthogonal quadratures of the electric field, known as the continuous variable (CV) approach Lloyd and Braunstein (1999); Fabre and Treps (2020). CV quantum photonics brings an advantage over DV approaches, namely the deterministic generation of entanglement over large photonic quantum states Weedbrook et al. (2012); Menicucci (2014). Millions of entangled optical modes have been experimentally generated in CV systems Asavanant et al. (2019). This makes CV photonics ideal for measurement-based QC (MBQC) implementations Menicucci et al. (2006), in which quantum computation is performed via single-qubit measurements on a large entangled quantum state. In photonics, single modes or collections of modes in the entangled state act as the analogues of qubits in matter-based approaches.
The key concept to enable quantum computation with large-scale entangled states, or in general with a large number of qubits, is that of fault tolerance. Fault-tolerant quantum computing (FTQC) is an approach to QC in which a computational (logical) qubit is encoded in many physical qubits on which quantum error correction (QEC) protocols are employed to correct for errors such as decoherence and photon loss Gottesman et al. (2001); Cai et al. (2021); Xu et al. (2023); If the probability of physical errors is below a certain value determined by the QEC protocol, called the fault tolerance threshold, the probability of logical error can be exponentially suppressed. There exist no direct fault-tolerant protocols for quadrature encoded continuous-variable quantum computing (CVQC) Ohliger (2010). An alternative approach is to discretise the CV space into qubits Ohliger (2010); Menicucci (2014).
The most promising qubits for CV FTQC are the Gottesman-Kitaev-Preskill (GKP) qubit Gottesman et al. (2001) and the magic states Bravyi and Kitaev (2005). Native logical gates are implemented with Gaussian operations on GKP qubits, while universal quantum computation is possible with the ability to create magic states. In fact, using only CV Gaussian resources combined with GKP states, one has all the necessary ingredients to achieve universal CVQC Baragiola et al. (2019). Therefore, much effort in CVQC has gone into devising scalable methods of producing high quality GKP qubits (that is, with high effective squeezing; see Sec. II), but producing these states with high success rates in photonics has remained a challenge (see Sec. V). After obtaining high quality GKP and magic qubits, CV architectures achieve fault tolerance by concatenating a second qubit-level error correcting code with the GKP states as the physical qubits. To implement these codes with optical fields, photonic switches were proposed either to increase the low success probabilities of GKP state generation Bourassa et al. (2021); Tzitrin et al. (2021) or for quantum error correction Larsen et al. (2021) (see Sec. VIII). Great progress has been made in lowering the noise and losses in photonic switches, but the loss levels are two orders-of-magnitude higher than passive photonics Team (2024), making removal of switches from the architecture a desirable goal.
With this context, we present four contributions in synthesizing an end-to-end CV FTQC architecture. First, we provide a protocol that produces fault-tolerant CV photonic qubits –GKP qubits– at state-of-the-art probabilities of more than $90\%$. Notably, the architecture makes use of all physical qubits, both above and below the FTQC threshold, to construct logical qubits. Second, we provide methods to generate GKP qubits on an integrated switchless photonic chip, opening up the possibility of low-loss manufacturability of the photonic chips with passive components. Our methods for state generation require only low photon number resolution (PNR) for which either superconducting microstrip single photon detectors (SMSPDs) Kong et al. (2024) or room temperature photon number resolving detectors Nehra et al. (2020) would suffice. Third, we present protocols for magic state generation with state-of-the-art success probability, an order of magnitude better than existing methods, for universal FTQC with feasible resource requirements. Fourth, we present the results of our quantum error correction simulations including GKP resource generation costs. We arrive at the gaussian squeezing threshold, the quantum noise metric in CVQC Menicucci (2014); Schnabel (2017), of $12$ dB to $13$ dB for fault tolerance, and distillable magic state success probability of $4.8\%$ also at $13$ dB. To our knowledge, the presented simulations are the first end-to-end CVQC simulations incorporating a realistic process producing GKP states with a range of effective squeezing values for fault-tolerant operation.
In the following sections, we start by providing an overview of photonic CVQC. We then describe our architecture before proceeding to discussing the details and physics of each architectural component, and conclude by providing the results of the fault tolerance simulations.
II CVQC Toolkit and Non-Gaussian States
In photonic CV MBQC, the set of entangling operations, non-destructive gates, and destructive measurements come from the CV photonics toolkit Pfister (2019), the most prominent components of which are the squeezing operations, linear optics, quadrature displacements, homodyne measurements, and photon number resolving measurements. In the following, we define the key terms, including entangled states, measurement operations, and non-Gaussian modes, which form the basic building blocks of our architecture.
The entangled state often used in photonic CV MBQC is the cluster state, canonically defined as Gaussian-squeezed modes entangled using controlled-phase gates ($C_{Z}$ielsen (2006). While CVQC was initially investigated with Gaussian cluster states requiring inline squeezing due to the use of controlled-phase gates Raussendorf and Briegel (2001); Flammia et al. (2009), an alternative formalism exists where inline squeezing is replaced by offline squeezing and passive interferometers Menicucci (2011); Walshe et al. (2021), resulting in a macronode cluster state, where collections of physical optical modes can be treated as a single mode on an underlying lattice Alexander and Menicucci (2016). An example of macronode cluster is the multimode state called a dual rail quantum wire, generated using the two-mode-squeezing operation implemented in a non-linear crystal followed by passive beamsplitters Pysher et al. (2009).
The name squeezing stems from the shape of the Wigner quasiprobability distribution of the conjugate photonic quadratures (denoted $p$ and $q$), in which the distribution for one quadrature has a narrower width compared to a coherent state, while the other quadrature has a broader distribution. The term also refers to the quantum noise on a measurement along one or a combination of the quadrature axes in phase space. In CV MBQC, because all gates are carried out via measurement operations, the noise on these gates is parametrized by the squeezing observed in each measurement. Thus, physical gate noise is directly tied to the squeezing level associated with each mode being measured. Ideal Gaussian squeezed states (infinitely squeezed) are denoted by |$\ket{m}_{p}$, where $m$ is the eigenvalue of the quadrature $p$. For the results presented in this paper, all simulations are run with finitely squeezed states (i.e., resulting from applying the finite squeezing (single-mode or two-mode) operator to vacuum). FT thresholds in terms of squeezing are important for CV FTQC as single and two-mode squeezing forms the fundamental resource for CVQC Gu et al. (2009); Menicucci et al. (2007).
The operations used to project a quantum state onto quadratures eigenstates are called homodyne measurements Lvovsky and Raymer (2009): a homodyne measurement with outcome $m$ on quadrature $p$ projects the measured mode onto the state |$\ket{m}_{p}$. An important property of homodyne measurements chosen with appropriate bases is that they can be used to either delete a node from a cluster state or teleport the quantum information from one node to another Gu et al. (2009), allowing one to shape cluster states during quantum computation.
While deterministic and scalable Gaussian entanglement generation gives us a substrate quantum state to use in CVQC, and homodyne measurements can be used to shape the cluster state and implement quantum gates, universal CVQC requires additional resources to enable quantum computations that are not efficiently classically simulable Bartlett et al. (2002). One prominent operation to enable universal CVQC is photon subtraction Ourjoumtsev et al. (2006), which can be used to implement gates consisting of arbitrary powers of the quadrature operators on squeezed states. Experimentally, in quantum optics, the most common method for performing photon subtraction involves using a highly reflective beamsplitter, with a photon detector—capable or not of resolving the number of detected photons— placed in the transmission arm Ourjoumtsev et al. (2006); Melalkia et al. (2022). Throughout this article, whenever the photon subtraction process is mentioned, we refer to this experimental setup. One of the key applications of these higher-order gates is the production of non-Gaussian states in an entangled cluster, which along with homodyne detections and displacements, give us the ability to apply quantum error correction and operate the universal quantum computer above the FT threshold Menicucci (2014).
The fundamental non-Gaussian resource states in our architecture are the squeezed cat states |±$\ket{\pm\alpha,r}$, which are superpositions of displaced squeezed states, i.e., |±$\ket{\pm\alpha,r}\sim(D(\alpha)+D(-\alpha))\mathcal{S}(r)\ket{0}$, where $D$ is the displacement operator in the phase space, $D(\alpha)=\exp{(\alpha a^{\dagger}-\alpha^{\star}a)}$, $\mathcal{S}$ is the squeezing operator, $\mathcal{S}(r)=\exp(\frac{1}{2}(r^{*}a^{2}-ra^{\dagger^{2}}))$, and $a^{\dagger}$ and $a$ are the creation and annihilation operators, respectively. Real $\alpha$ results in $q$ quadrature displacements ($D_{q}$), while imaginary values result in $p$ displacements ($D_{p}$). Cat states are parametrized by squeezing ($r$) and displacement ($\alpha$), which we take to be real-valued in the following results. The Wigner function for a squeezed cat state is shown in Fig. 1. While FTQC codes exist with cat states as the logical qubits, these require the experimentally-challenging parity check operator Schlegel et al. (2022), limiting the role of cat states in photonic FTQC to that of resource states which we consume to produce qubits.
GKP states, and more generally grid states, were first proposed in [Ref. Gottesman et al. (2001)], with a considerable amount of effort spent studying their use in fault-tolerant architectures in recent years Tzitrin et al. (2020); Mensen et al. (2021); Grimsmo and Puri (2021); a typical Wigner function is shown in 1. The sensor state, denoted |$\ket{S}$, is a symmetric grid state defined as the $+1$ eigenstate of the displacements $D_{q}(\sqrt{2\pi})$ and $D_{p}(\sqrt{2\pi})$, while a GKP qubit is the $+1$ eigenstate of the displacements $D_{q}(2\sqrt{\pi})$ and $D_{p}(\sqrt{\pi})$, where we have set $[a,a^{\dagger}]=1$  Gottesman et al. (2001); Grimsmo and Puri (2021) (see Appendix Sec. XI.1 for more details). Interference between sensor states using passive linear optics can be used to produce entangled GKP Bell pairs  Walshe et al. (2020).
Ideal GKP states have wavefunctions that are a series of delta functions in the $q$-basis (or $p$-basis) of the continuous-variable $(q,p)$ phase space of a bosonic mode, and are therefore unphysical (see Appendix XI.1 for ideal GKP wavefunctions). Physically-relevant approximations to the ideal GKP states can be obtained by applying the damping operator to the ideal state Matsuura et al. (2020). The width of the peaks provides an effective squeezing parameter, which describes the closeness of a given GKP state to the ideal state. The probability of physical gate errors in CVQC depends on the effective squeezing of the underlying GKP states, which denotes the noise level of measurements associated with the qubit compared to the standard noise of a coherent state Weigand and Terhal (2018); minimizing this noise corresponds to reduced physical error rates. For any given density matrix $\rho$ and the desired displacement operator with displacement $u$, we estimate the effective squeezing of the output state as $\Delta_{D}=\frac{1}{u}\sqrt{\ln{|TrD(u)\rho|}}$ where $Tr$ is the trace operation Weigand and Terhal (2018). The quality of a realistic GKP state for FTQC is quantified most readily by the effective squeezing parameter since the probability of error in the logical X or Z operations is bounded by $P<\frac{2\Delta}{\pi}\exp{-\pi/(4\Delta^{2})}$, where $\Delta$ is the effective squeezing in the respective quadrature Weigand and Terhal (2018). Note that the GKP state can have different effective squeezing in the two quadratures owing to different parameters at play in the generation protocols. A final important non-Gaussian state in our architecture is the magic state. These are essential for universal quantum computation by providing an implementation of a complete gate set. In CVQC, magic states are also grid states; their Wigner functions contain periodic structure (see Fig. 1; also Appendix Sec. XI.1), and they can be derived from CV GKP states Baragiola et al. (2019). An important metric is the distillation threshold of magic states, which quantifies the fidelity above which many copies of noisy magic states can be used to distill a high fidelity magic state useful for applying fault-tolerant non-Clifford gates, e.g., the $T$ gate Bravyi and Kitaev (2005).
Figure 1: Wigner functions of squeezed cat state, GKP zero qubit |$\ket{0}_{G}$, and magic H state |$\ket{H}$. The horizontal and vertical axes are the $q$ and $p$ quadrature axes, respectively. Each of the states has squeezing of approximately 13dB in the $q$ direction (cat state) or both the $q$ and $p$ directions (GKP and magic states). Red peaks correspond to negative values of the Wigner distribution, while blue peaks are positive. The color scale is otherwise arbitrary and is meant only to notionally represent the location of probability density.
III Architecture and Design Overview
Our photonic QC architecture contains three key components for physical qubit generation (see Fig 2): cat state generation via the Photon-counting-Assisted Node-Teleportation Method (PhANTM) Eaton et al. (2022a) (see also section IV), an adaptive breeding protocol (detailed in section V), and a magic state generation protocol (described in section VI). The logical layer consists of two major components: logical qubit generation, which consists of distributing physical qubits as the nodes within the Raussendorf-Harrington-Goyal (RHG) lattice Raussendorf et al. (2006); and the quantum error correction (QEC) encoding/decoding system, which is also responsible for running logical instructions. All of the physical qubit generation and RHG lattice generation components are possible on a set of integrated photonic chips without active photonic switches or variable beamsplitters on the quantum modes to minimize loss and noise. While fast electro-optic modulation is needed on local oscillators during homodyne measurements, the local oscillator power can be corrected to account for any losses.
We use cluster engineering with time multiplexing and photon-number-resolving detection to prepare cat states, followed by adaptive breeding to create GKP states using homodyne measurements. Magic states are produced by applying CV QEC on cat states, using bred GKP states as resource states. We generate the RHG lattice using static linear optics and perform syndrome extraction and logical operations by single qubit homodyne measurements on logical qubits within the surface code. Classical compute resources are used to implement the user interface, code compilation, and transpilation to local oscillator phase controls, which implement all gate controls in the architecture. Homodyne detection measurement results are sent to the logical instruction and QEC system where syndrome decoding and error tracking are performed. Any necessary modifications to the measurement results as a consequence of error correction are applied before using measurement outcomes to compute the algorithm results. When online changes to future measurements are required, the result of error detection is sent to the phase modulators which control the local oscillator phases that set the measurement bases.
Figure 2: Conceptual overview of the creation of GKP states, magic states, MBQC-based lattices, and QEC to enable FTQC. Repeating photon subtraction and teleportation along a cluster state generates cat states with a high success probability and low photon number resolution. The cats are then squeezed to produce cats with deterministic amplitudes which are bred together to create GKP states without the need for post-selected measurements. All steps happen on the cluster state. The GKP states and cat states are used together to create magic states. The architecture produces $H$-type magic states in this configuration. Notations used in circuits are defined in main text. Additionally, $\mathcal{S}$ and $R$ stand respectively for squeezing and rotation operators. Note that for brevity we use the notation for infinitely squeezed Gaussian states in the circuits, but in practice they are replaced by finitely squeezed states in the simulations. Downward arrows in circuits are beamsplitters while vertical lines with circular caps are $C_{Z}$ gates.
The substrate to create GKP and magic states is a cluster state formed from two-mode squeezed light and passive beamsplitters, where each node is a photonic mode characterized by the quadratures $q$ and $p$, and features nearest-neighbor entanglement between the nodes. The time-multiplexed cluster state in our architecture is indexed by spatial and optical frequencies along with temporal modes (see Fig. 4 and later figures on cluster state engineering). To produce cat states our architecture uses PNR detection at specific optical frequencies and timesteps. Specifically, Schrödinger cat states are produced using the Photon-counting-Assisted Node- Teleportation Method (PhANTM) Eaton et al. (2022a). PhANTM is well-suited to our time-frequency-space multiplexed architecture, as individual nodes of a cluster state can be made non-Gaussian with photon subtraction, and the non-Gaussianity can be teleported through a cluster state with homodyne measurements. This allows us to repeat the PhANTM process over multiple time steps to increase the amplitude of the generated cats and make them suitable for high-probability GKP generation, in contrast to single-shot GKP generation methods without time or frequency multiplexing Bourassa et al. (2021); Takase et al. (2024).
The cat states, which are embedded in the cluster state, are squeezed using teleportation-based squeezing gates Alexander et al. (2014) and then “bred” into either sensor states for qubit production or GKP states for magic state production Baragiola et al. (2019). Switching between the two output states is achieved by changing the required amplitude of the input cat states. Given that our protocols produce cat states with probabilistic squeezing and amplitude located at the nodes of a cluster state, we use an adaptive breeding protocol to produce GKP qubits or sensor states. In our adaptive protocol, cat states of various multiples of the desired grid state spacing are used including a single mode squeezed vacuum that can be substituted for lower quality cat states when desired (see Algorithm  V).
These fundamental components can all be implemented in passive photonic integrated circuits, with squeezed light, sensor states, magic states, and the RHG lattice all produced on integrated chips. The chips integrate linear optics components, such as beamsplitters and phase shifters, and nonlinear components such as $\chi^{(3)}$ media, as shown in Fig. 3. A squeezing resonator is bidirectionally pumped by pumps centered at $\nu_{0}$ with $\nu_{p_{1}}=\nu_{0}+\Delta\nu/2$ and $\nu_{p_{2}}=\nu_{0}-\Delta\nu/2$, where $\Delta\nu$ is the free spectral range (FSR) of the resonator. This creates two squeezed combs each centered about the respective pump. Interfering the two squeezed combs at a balanced beamsplitter results in the formation of a 1D dual-rail quantum wire Alexander et al. (2016). Subsequently, homodyne measurement are performed locally within this cluster to reshape it in preparation to apply PhANTM or adaptive breeding at different time steps. If the local oscillator phase is aligned with the $q$ quadrature then the node is removed whereas if the phase is that of the $p$ quadrature, the quantum information is teleported to a neighboring node, with an additional $\pi/2$ rotation applied Menicucci et al. (2006). Each frequency mode needs to be separated from the optical frequency comb using frequency filters to then be detected. These filters are schematically shown as ring resonators, which are among the most compact and lowest-loss options for frequency separation Liu et al. (2021) negligibly impacting the pass-through frequency modes, and demonstrating very high transmission efficiencies of the filtered mode. The delay is used for time multiplexing before a final balanced beamsplitter is used to interfere frequency modes in (b) with the delayed modes in (a) to produce the final cluster state. With this resource state CVQC operations can be applied to generate cat states by performing PhANTM, squeezed cats using a squeezing gate, or GKP states by adaptive breeding.
Figure 3: Chip design of GKP generator to create multiple cat states in parallel from a time-frequency encoded Gaussian cluster state. The required squeezing is produced via four wave mixing in a ring resonator ($\chi^{3}$). Cat states are bred within the same chip to generate GKP state. The shape of the cluster state can be adapted as function of the operation applied in stage IV. $u_{p1}$ and $u_{p2}$ stand for the optical pump beams at frequency respectively $\nu_{p1}$ and $\nu_{p1}$. Boxes with $p/q$ measurement indicate circuits where homodyne detection and corrective displacement are performed while boxes denoted $\mathcal{O}$ represent circuit where multiple photons subtraction are applied as well as displacement and homodyne detection. The green circles denote frequency filters that transmit given frequency modes. The labels within the circles correspond to the frequencies mode number consistent with those indicated in Fig. 4.
The architecture implements quantum error correction (QEC) by embedding the bred GKP sensor states from the photonic integrated circuit in Fig. 3 into the RHG cluster state Raussendorf et al. (2006), the MBQC analog of the surface code. Large scale entanglement is generated using a macronode based approach Tzitrin et al. (2021) where each qubit of the RHG lattice is replaced by four optical modes. Entanglement links in the RHG lattice are replaced by GKP Bell pairs and the final cluster state is then generated by interfering all modes of the macronode with passive linear optics.
Logical gates can be implemented on the RHG cluster state using lattice surgery Herr et al. (2018) and magic states. Magic states can be embedded in the macronode RHG lattice by replacing a GKP Bell pair with the result of interfering a magic state with squeezed vacuum Tzitrin et al. (2021). More detail on the lattice construction is given in Sec. VII.
IV Cat States on dual-rail quantum wire
Cat states are a fundamental resource in our CVQC architecture. They are used as a resource state to generate both sensor states via breeding and magic states via implementing CV-QEC on input cat states. Methods for creating cat states in CV photonics need at a minimum a combination of Gaussian and non-Gaussian resource states, or photon number detection along with Gaussian resource states. A method was proposed to create cat states by applying PNR on squeezed vacuum using machine learning to control some of the experimental parameters Anteneh et al. (2024). This method is efficient for state generation, requiring only a few iterations of photon detection, but requires high photon number resolution. Another proposal used high-photon-number Fock states (with $10$ or more photons) along with photon number resolution to create cat states deterministically Winnel et al. (2024). Creating high photon-number Fock states presents a high experimental challenge Uria et al. (2020). In PhANTM Eaton et al. (2022a), the method used in the present architecture, applications of probabilistic photon subtraction and teleportation build up non-Gaussianity over time. This requires a Gaussian cluster as a resource to repeatedly apply the method. One of the key advantages of PhANTM lies in its ability to generate sufficiently large cat states that can subsequently be bred while distributing the total number of photons subtracted over multiple repetitions of the operation.
Figure 4: Cluster preparation for PhANTM application and parallel dual-rails. I, II, III and IV refer to locations on the chip in Fig. 3. In stages I to III, the cluster state is generated and then reshaped using homodyne measurements to form dual rails independent of each other. In IV, PhANTM is applied as the circuit in Fig. 5 shows: at the time step $\tau_{i}$ modes $\nu_{a}$ and $\nu_{b}$ are respectively mode 2 and 1 in the circuit. Then the output mode in the dual rail is the mode $\nu_{b}$ of the $\tau_{i+1}$ time step. Dark blue and maroon segment represent effective $C_{Z}$ gates with weights of respectively $-1$ and $+1$, coming from applications of two-mode-squeezing and beamsplitter operations, while red circles represent the second beamsplitters in Fig. 3, where a phase can be applied to correct any rotation induced by cluster engineering. Orange and blue segments represent $C_{Z}$ gates with respective weights of $+1/2$ and $-1/2$. The red arrows in II show the direction of the teleportation.
Fig. 4 shows the steps and cluster engineering using homodyne measurements, leading up to many copies of cats generated in parallel from a single frequency-space-time entangled cluster state. The frequencies are denoted by $\nu$ or arabic numerals, spatial rails are denoted by (a) and (b), and time steps are denoted by $\tau_{i}$. Starting with many pairs of two-mode entangled states in I interfered at the first beamsplitter of Fig. 3, we arrive at the entangled frequency chain also known as the dual-rail quantum wire shown in Fig. 4 II. We apply $p$ and $q$ homodyne measurements in region II of Fig. 3 and Fig. 4 to convert a dual rail wire into multiple wires in frequency-space as shown in region III of Fig. 4. Going through a time delay then entangles the two-mode wires in time to create many dual-rail wires in frequency-space-time as shown in Fig. 4, on which PhANTM can be applied in parallel, e.g., on modes $2_{a,b}$ and $10_{a,b}$ in Fig. 4 at time $\tau_{i}$. Since the cluster state modes that undergo homodyne detections or PhANTM are pre-determined, no active switching is required.
The application of PhANTM on the dual-rail wire in Fig. 4 IV corresponds to repeating the circuit in Fig. 5 in time. The rotation on the output mode can be manipulated through the phase of the beamsplitter and homodyne detection basis. The beamsplitter is defined by two parameters, $\theta$ and $\phi$, which are related to the reflectivity and transmitivity such that $r=e^{i\phi}\sin(\theta)$ and $t=\cos(\theta)$. We have chosen the phase $\phi=\pi/2$ in Fig. 5 such that a single application of the circuit at $\tau_{i}$ produces an output cat with the same orientation as the input cat, allowing repeated applications of PhANTM at $\tau_{i+1}$ and later timesteps. Note that there are two homodyne measurements applied in a unit round of PhANTM (one on $\nu_{a}$ and another on $\nu_{b}$) before the output mode is ready for the next round. Also note that given the beamsplitters and phases involved in the circuit in Fig. 5, the bases used for either mode teleportation or deletion differ from the usual $p$ or $q$ bases for a canonical cluster state  Raussendorf (2003). The effective $C_{Z}$ gate in Fig. 5 results from the application of two-mode squeezing, beamsplitters, and homodyne measurements as shown in Fig. 4. The modes that undergo operation $\mathcal{O}$ in Fig. 3 have multiple photon subtractions followed by homodyne detection applied to them, while the rest of the modes in IV have $p$ basis homodyne measurement applied to them. Active displacements are needed in between repeated applications of PhANTM due to random homodyne measurement results. The value of the displacement is the output of the homodyne detector fed directly back to a standard displacement circuit with a local oscillator for the next temporal mode Eaton (2022). The timescale of these active displacements must be shorter than the time delay between successive temporal modes in the cluster state.
Figure 5: Application of PhANTM at stage IV. The top two modes are entangled through a beamsplitter (downward arrow) while the bottom two are entangled through a $C_{Z}$ gate (a line with two nodes at the ends) which is implemented by two-mode squeezing, beamsplitter, and homodyne measurements for cluster engineering. Conceptually, the $C_{Z}$ gate happens first, followed by the beamsplitter, and then the measurements from top to bottom. $O$ stands for multiple photon subtraction operations.
Monte Carlo simulation of several rounds of PhANTM were conducted using the unit cell shown in Fig. 5, sampling the number of photons detected in each photon subtraction. The parameters of the cat states that we aim to optimize are their size and squeezing, both of which are stochastic. The larger these parameters, the higher the effective squeezing of the resulting GKP states, leading to lower error rates for QEC codes. To facilitate analysis and comparison across various conditions, a single parameter is defined: the amplitude of the cat state with zero squeezing ($\alpha_{c}$, also see Sec. V). Fig. 6 shows the mean of $\alpha_{c}$ as a function of cluster squeezing for different number of PhANTM rounds (see also Sec. XI.3 in Appendix). In these simulations, eight sequential photon subtraction attempts were performed at each time step. The beamsplitter angles for each photon subtraction are chosen such that no more than $10$ photons are detected per PNR detector. The subtraction probability goes down for later subtraction events, as continued subtraction attempts de-amplify the quantum signal. We apply a reflectivity gradient (higher angle in later subtraction attempts) to ensure that the distribution of subtracted photons among the PNR detectors is roughly uniform (see Sec. XI.4 in Appendix). Increasing the cluster squeezing enhances the potential number of photons that can be subtracted in a single subtraction event, resulting in cats with larger amplitudes. Furthermore, increasing the number of PhANTM steps raises the total number of photons subtracted across all cumulative PhANTM events, leading to larger amplitudes for cat states. The additional squeezing value also depends on the reflection and transmission coefficients of each beamsplitter utilized in the photon subtraction process.
The success probability of each step of PhANTM, and the final cat size, can be tuned given the hardware constraints of the photon number resolution ceiling and the losses in the photonic circuit resulting from repeated teleportations in time, leading to many possible optimizations.
Figure 6: Mean $\alpha_{c}$ as a function of cluster squeezing. Each color corresponds to a given number of PhANTM steps. The means are estimated from 500 Monte Carlo iterations of PhANTM. Details of these simulation are given in Appendix. The plot shows the positive correlation of $\alpha_{c}$ with cluster squeezing and the total number of PhANTM steps.
V Generation of GKP States
[…]
We focus on creating GKP states above the FT threshold from breeding a supply of probabilistic cat states from PhANTM. Since PhANTM is a probabilistic process, fixed-$\alpha$ and fixed-parity cats are difficult to engineer. In addition, with a low probability the cat states produced from PhANTM have amplitudes that are too small for breeding. Finally, given a finite amount of squeezing in the cluster state, some cats from PhANTM are too large to be squeezed to the amplitude required for breeding. In the following, we present an adaptive breeding protocol that overcomes the probabilistic nature of PhANTM to create GKP states with a high effective squeezing.
The pseudoalgorithm for adaptive breeding is shown in Algorithm V. The algorithm takes a set of cat states embedded in a cluster state. In order to breed GKP or sensor states with the correct spacing, the cat amplitudes must be adjusted using a squeezing operator. Once the amplitudes are adjusted, the cats are simply measured in the $p$ basis. Because the cats are embedded in a cluster state, they have effective $C_{Z}$ operations applied between them, and squeezing and breeding is thus implemented via measurements only Eaton et al. (2022a).
{algorithm}
GKP adaptive breeding algorithm Input cluster squeezing $r_{cluster}$
Input $M$ breeding steps
Input $2^{M}$ outputs from PhANTM, each with: unsqueezed cat amplitude $\alpha$, parity $P$, cat squeezing $r_{cat}$
Input desired GKP state spacing $\xi$
Input desired cat $\alpha_{b}$ in Weigand and Terhal (2018) $\alpha_{b}=\xi 2^{(M-3)/2}$
Input lower limit of acceptable cat squeezing $r_{lb}$
Input measurement bases $p$ and $q$
for each PhANTM output do
     if $P$ is odd then
         displace cat to change phase by $\pi/2$
     end if
     compute $s_{\alpha_{b}}=\ln(\alpha_{b}/\alpha)$ and $s_{2\alpha_{b}}=\ln(2\alpha_{b}/\alpha)$
     if $s_{\alpha_{b}}<r_{cluster}$ and $s_{\alpha_{b}}>r_{lb}$ then
         squeeze cat with $s_{\alpha_{b}}-r_{cat}$
         teleport with $p$
     else if $s_{2\alpha_{b}}<r_{cluster}$ and $s_{2\alpha_{b}}>r_{lb}$ then
         squeeze cat with $s_{2\alpha_{b}}-r_{cat}$
         teleport with $p$
     else
         measure with $q$
     end if
end for
begin Breeding Procedure:
for each output pair produced above do
     select unmeasured modes on the cluster state on which to teleport output bred states
     breed and entangle outputs by measuring one mode of pair with $p$
end for
for each unmeasured mode remaining above do
     select unmeasured modes on the cluster state on which to teleport output bred states
     breed and entangle outputs by measuring one mode of pair with $p$
end for
repeat Breeding Procedure until only one mode remains unmeasured
The adaptive breeding protocol takes several inputs: the number of desired breeding rounds, $M$; the squeezing present in the cluster state, $r_{cluster}$; a number of input cats which depends on the number of breeding rounds, $2^{M}$; the input cats’ unsqueezed amplitudes $\alpha$, their present squeezing level after PhANTM, $r_{cat}$, and their parity $P$; the desired GKP or sensor state spacing in phase space, $\xi$; the required cat state amplitude for breeding, $\alpha_{b}$, which depends on the spacing as $\alpha_{b}=\xi 2^{(M-3)/2}$ Weigand and Terhal (2018); and the lower bound on cat squeezing that will determine its suitability for breeding, $r_{lb}$. All measurements are performed in either the $p$ or $q$ basis depending on whether an input cat is used in breeding or replaced with a squeezed vacuum.
Figure 7: Cluster engineering for adaptive breeding. While the figure follows the same structure as in Fig. 4, a notable difference is the connection between dual rails of different frequency modes required here. Breeding is applied between states in modes $4_{a}$ and $8_{a}$ at the time step $\tau_{i+2}$ in IV. The output state of the breeding protocol is in mode $6_{a}$. Also at stage IV and prior to breeding, squeezing gates are applied at time step $\tau_{i+1}$ by setting angles of the homodyne detections. Note that the physical components on the chip create the cluster state in I and II, which is then engineered for different processes in III and IV by changing homodyne measurements.
As a first requirement for the adaptive breeding to proceed, the amplitude $\alpha$ of the cats output from PhANTM must be known. The output amplitude and squeezing, $r_{cat}$ depends on the photon subtraction vector obtained from PhANTM. The parity $P$ of each cat also depends on the total number of photons subtracted during PhANTM. As each cat can only breed with another cat of the same parity, small displacements are used to adjust the odd cat parities to even ones. This displacement causes an effective, known displacement in the first breeding round output, which can be kept track of but does not have to be actively fixed until breeding is complete. In general, the cat amplitudes are adjusted via squeezing operations of magnitude $s_{\alpha_{b}}$, which depends on the desired spacing $\xi$ and the number of breeding steps $M$.
The extra teleportation step in the breeding protocol allows us to introduce momentum squeezed states at $r_{cluster}$ squeezing in two scenarios. If the cat amplitude produced by PhANTM is too small to be used in GKP generation such that no amount squeezing smaller than $r_{cluster}$ can be used to adjust its amplitude to $\alpha_{b}$, it can be replaced by a momentum squeezed state before breeding. Adding a squeezed vacuum state in place of a cat increases the squeezing in one quadrature at the expense of the squeezing in the other quadrature. In the second scenario, arbitrarily large cat states can be produced in PhANTM that are too big to be squeezed to amplitude $\alpha_{b}$ needed in breeding because the squeezing required to do so would exceed $r_{cluster}$. These cats can be squeezed to an even multiple of the required amplitude for the breeding protocol to proceed ($2\alpha_{b}$ is used in Algorithm V). If the squeezing of these large cats, $s_{2\alpha_{b}}$, is smaller than the desired lower bound, $r_{lb}$, then these cats can once again be replaced by momentum squeezed states (and likewise if the required squeezing would exceed $r_{cluster}$). The squeezing of cats or cluster is the dominant term that determines the effective squeezing of the GKP states in one quadrature, while the number of breeding rounds is the dominant term for effective squeezing in the other direction. Note that using a cat that is an even multiple of the required cat squeezing also has the effect of increasing the effective squeezing in one of the quadratures, giving us flexibility in trading off the squeezing in the GKP quadratures by combining mixtures of cats with different amplitudes and momentum squeezed states.
Fig. 7 illustrates the cluster engineering method required to apply adaptive breeding in Algorithm V to cat states generated on different dual rails of the cluster state (see also Appendix Sec. XI.5). At stage II, the modes arrangement is modified compared with Fig. 4. Specifically, prior to applying any breeding operation, distinct frequency modes need to be connected via an effective $C_{Z}$ interaction. This connection is shown in stage III. Breeding of cat states in mode $4$ and $8$ is applied in stage IV at the time step $\tau_{i+2}$. Prior to adaptive breeding, squeezing gates are employed to ensure cat states sizes match $\alpha_{b}$ or an even multiple of $\alpha_{b}$ in the pseudoalgorithm above. This operation requires one single time step and consumes two spatial modes of the same frequency as it is represented between time step $\tau_{i+1}$ and $\tau_{i+2}$ in Fig. 7. The squeezing parameter of the gate is set by the homodyne measurement angles $\theta_{a}$ and $\theta_{b}$ Alexander et al. (2014). More details are discussed in Sec. XI.5 of the Appendix. Note that we can swap between sensor states and GKP qubits by changing the amplitudes to which the cats are squeezed.
The outputs of the previously described PhANTM simulations are used as input to an adaptive breeding simulation. A Monte Carlo approach is used to obtain statistics about the sensor states obtained from adaptive breeding, with 1000 samples drawn from a set of input cats that were produced with an array of settings for starting cluster squeezing and the number of photon subtraction attempts. The simulation makes use of Qutip Johansson et al. (2013) for density matrix calculations and Strawberry Fields Killoran et al. (2019) for homodyne detection simulation. The average sensor state squeezing along with their standard deviations are shown alongside FT threshold simulations in Sec. VIII.
VI Generation of magic States
Magic states are an important ingredient of universal FTQC as they allow implementation of non-Clifford operations while themselves remaining distillable with only Clifford gates Bravyi and Kitaev (2005). A popular gate set which generates the Clifford group is {$CNOT$, $H$, $S$}, where $CNOT$ is the controlled NOT gate; $H$ is the Hadamard gate, and $S$ is the phase gate. The non-Clifford T gate usually completes the universal gate set Bravyi and Kitaev (2005). Provided with a source of physical magic states, one can distill a resource state that enables teleportation of non-Clifford operations into the QEC substrate at arbitrarily low error rates. Improving magic state quality by increasing the fidelity with respect to any of the |$\ket{H}$ or |$\ket{T}$ states reduces the physical resources required for a given logical error rate in the quantum computation. A promising method to produce a GKP magic state uses GKP error correction on the vacuum state Baragiola et al. (2019), implying that GKP qubits and encoded Clifford operations are sufficient for universal QC Baragiola et al. (2019); Calcluth et al. (2022). However, simulations show that a reasonable success rate with existing methods requires the effective squeezing of GKP states to be extraordinarily high, approaching $20$ dB (see Fig. 8), to create distillable magic states in one error correction round. In other architectures, producing magic states with practical success rates require the GKP states to have very high effective squeezing Baragiola et al. (2019); Takase et al. (2024), well above the FT threshold. In this section we provide a method to circumvent this issue and produce magic states with GKP resources that have effective squeezing levels near the FT threshold.
The method used to create magic states with significant probability using much lower squeezed GKP states in the present architecture is shown in Fig. 9. We replace the input vacuum with a cat state (|±$\ket{\pm\alpha_{4}}$ in Fig. 9) on which to apply QEC with GKP states. This substitution opens up the possibility to optimize the input cat amplitude, as well as all operations in the GKP breeding process prior to CV-QEC. We also replace the $C_{Z}$ operations during CV-QEC with beamsplitters, which allows us to use separate photonic chip components to produce GKP and cat states.
Figure 8: Top: magic state infidelity as a function of squeezing in the GKP states used as a QEC resource for three input cat amplitudes ($\alpha$), where $\alpha=0$ corresponds to the vacuum state. Bottom: magic state infidelity as a function of input cat amplitude and GKP squeezing of 13 dB. The horizontal line represents the distillation threshold for H-type states.
This substitution requires a single squeezing operation to be performed on one of the GKP resource states, which is easily commuted to the state preparation stage before breeding the GKP state. A simple implementation of the protocol is shown in Fig. 9, resulting in a magic state generation circuit using nine cat states. This contains the equivalent of two GKP breeding rounds for each of the two GKP resource states that are consumed; the circuit with 3 GKP breeding rounds would contain 17 input cats. The cat input |±$\ket{\pm\alpha_{4}}$ breeds successively with the output of two cat trees, each of which produces high quality GKP states.
The distillation infidelity threshold for H-type states is approximately 14.7%. The plots in Fig. 8 show a deterministic simulation of CV-QEC on input vacuum or cat states with post-selection on quadrature measurement outcomes of zero. Note that this post-selection, done to ease simulations, will slightly underestimate the output squeezing of the bred GKP states Weigand and Terhal (2018). We see that when performing CV-QEC on the vacuum mode (purple data points) one obtains infidelities above the distillation threshold (of 14.7%) when using GKP states with squeezing of 13dB or less as the resource state for CV-QEC. On the other hand, with effective squeezing greater than 12dB and cat states with amplitude of $\sim 5.1$, magic states with infidelities below the distillation threshold can be obtained.
Figure 9: Magic state generation protocol. Double lines represent feed-forward operations. The gauge boxes represent $p$ basis measurements, except for the final measurement which is in the $q$ basis. The unsqueezed input cats labeled |±$\ket{\pm\alpha_{i}}$ can also be replaced with prior cascaded breeding steps in order to increase the quality of the bred GKP states before applying CV-QEC.
In practice, the cat states produced by PhANTM have random amplitudes, and they must be squeezed in order to obtain the correct spacing for GKP breeding before magic state generation. This means that our GKP states have a random squeezing distribution. Further, as Fig. 8 shows, the optimal input cat amplitude for the error-corrected mode is between 4.9 and 5.5. Because our cats have random amplitudes and squeezing levels, only a few will hit this required target, although PhANTM can be engineered to emit cat states with mean squeezing levels and amplitudes centered on the requirements. Monte Carlo simulations were used to determine the success rate of producing a magic state when random cats from PhANTM were used to provide the resource states, both for GKP breeding and the cat state upon which CV QEC was performed. The simulations use Qutip Johansson et al. (2013) for density matrix calculations and Strawberry Fields Killoran et al. (2019) for homodyne detection simulation. Fig. 10 shows histograms of the infidelities obtained from each magic state generation attempt, where the fidelity is measured with respect to an ideal $H$-type magic state. The cat states were produced using Monte Carlo PhANTM simulations with a cluster state squeezing of 13.5 dB, 20 PhANTM steps, and 8 photon-subtraction attempts made before homodyne detection and teleportation to the next PhANTM time step. When using vacuum states as the error-corrected mode, we see a success rate of 0.003 after 1000 iterations (brown counts in the histogram). On the other hand, when using cat states as the error-corrected mode, we see a success rate of 0.048 after 1000 magic state generation attempts, an improvement of more than factor of 10, which reduces resources required to produce a magic state by a factor 10 relative to using vacuum alone. This simulation specifically targeted $H$-type states, which enable teleportation of $T$ gates within the architecture. These success rates mean that magic factories will not require the majority of qubit resources in surface code substrates using lattice surgery Litinski (2019a).
Figure 10: Histograms of fidelities obtained from Monte Carlo simulation of magic state generation for CV QEC on vacuum modes (brown) or cat states (red).
VII Encoding Logical Qubits in a Surface Code
GKP encoding can protect against small displacement errors, but larger displacements ($>\frac{\pi}{2}$) will result in an error. Depending on the quadrature, this can be either a Pauli X error ($q$ quadrature) or a Pauli Z error ($p$ quadrature). These logical qubit errors require additional qubit-level error-correction. Many error correcting codes have been studied in the context of GKP qubits Bourassa et al. (2021); Stafford and Menicucci (2023); Raveendran et al. (2022); Zhang et al. (2021). We will focus on topological error correction based on a family of codes known as surface codes, which have been shown to have high thresholds ($\approx 1\%$ under circuit level noise Wang et al. (2011) or $\approx 10\mathrm{dB}$ with GKP states Bourassa et al. (2021)) and well-studied logical gates Litinski (2019b).
As mentioned earlier, photonic CVQC is well-suited to MBQC. Circuit model codes known as Calderbank-Steane-Short(CSS) codes, where X and Z errors are detected independently, can be translated into a cluster state for fault-tolerant MBQC via a process known as foliation Bolt et al. (2016), which loosely maps the data qubits in the circuit model code to linear cluster states. Surface codes are examples of CSS codes, and the resulting foliated 3D cluster state is often called the Raussendorf-Harrington-Goyal (RHG) lattice after the authors who first introduced it Raussendorf et al. (2007). The unit cell of this cluster state is shown in Fig. 11. The product of $X$-basis measurements (equivalently $p$-basis homodyne measurements for GKP qubits) on all face qubits is a stabilizer of the cluster state. In the error free case, measuring stabilizers will result in +1 eigenvalue and errors will result in $-1$ measurement results for any stabilizer that they anti-commute with. A second unit cell can be defined by shifting the base unit cell by 1/2 in all directions. Here, the edge qubits of the initial unit cell become face qubits of the second unit cell. These two lattices are known as primal and dual lattices, respectively. By measuring the stabilizers of both lattices, errors on all qubits can be detected. Logical operators in foliated cluster states are formed by correlation surfaces that map logical operators from the underlying code from one end of the cluster state to the other. To implement logical X and Z operations, correlation surfaces of the RHG lattice are given by 2D sheets of qubits which span the lattice and intersect primal and dual unit cells, respectively. The code distance is the shortest chain of errors that can result in a logical error, which is equivalent to the shortest error chain that can span the RHG lattice. The threshold theorem states that provided the physical error rate of the underlying qubits is below some threshold, we can arbitrarily suppress the logical error rate by increasing the code distance Knill et al. (1998).
Figure 11: (a) RHG unit cells. Primal unit cell shown with blue qubits on faces and red qubits on edges. Products of $X$-basis measurement on face qubits are a stabilizer of the cluster state. Dual unit cell shown shifted by half a unit in all directions. Now red qubits are on faces and blue qubits are on edges. By measuring all primal and dual stabilizers, Pauli errors on all qubits can be corrected. (b) Six macronode tilable unit cells shown relative to the RHG unit cell. Purple arrows show where GKP nodes are sent and green arrows dictate delayed modes. (c) Macronode layout within the unit cell. 12 GKP Bell pairs are created per unit cell with four modes being sent to neighboring unit cells and four being received. Temporal connections between macronodes are created using delay lines. The colored links represent fiber optic connections between separate optical chips for the GKP generation and macronode interferometer circuits. (d) Example of the tileability of the unit cell into a larger RHG lattice. Orange links denote Bell pairs shared between unit cells and green edges indicate temporal links. Bell pairs can be generated by interfering sensor states output from the GKP generation algorithm on a beamsplitter.
The popular Clifford + T gate set maps naturally to the surface code. Clifford gates and teleportation based T gates can be implemented by lattice surgery Litinski (2019b). This approach defines logical qubits as patches, which are subsections of the overall surface code lattice, with logical operators identified on the edges. In MBQC, this corresponds to volumes of the RHG lattice with correlation surfaces on the boundaries. Patches can be merged by measuring stabilizer checks between edges of the patches, thus performing a joint logical measurement of the two patches. Clifford gates and the T gate can be implemented with joint measurements including ancilla qubits or magic states. In fact, in Ref. Litinski (2019b), it was shown how all Clifford gates can be commuted through the T gates and measurements, resulting in circuits that just consist of mulit-qubit joint measurements including a magic state. Lattice surgery operations can also be mapped to changing measurement bases (homodyne angles) in fault-tolerant MBQC Herr et al. (2018).
This gate set is well studied and optimized, however our architecture can be adapted to other gate sets Litinski (2019b); Gidney and Fowler (2019) and other error correcting codes Walshe et al. (2024). In this paper we aim to quantify Gaussian squeezing resources needed to demonstrate FT thresholds with the RHG lattice serving as the example code. We leave optimization of the error correcting code and gate set for future work.
VII.1 Macronode GKP lattice construction
With GKP qubits there exists more than one method for constructing the RHG lattice in Fig 11(a).
We can construct the lattice directly, where the nodes are GKP $\ket{+}$ states and the edges are CV $C_{Z}$ gates, which can be implemented deterministically with linear optics and squeezing. This approach has two main draw backs. Firstly, while it has been demonstrated previously Yoshikawa et al. (2008), in-line squeezing is generally difficult to realize experimentally. Second, ideal $C_{Z}$ gates are not physical, requiring infinite squeezing. Approximating the gate with finite squeezing will add noise to the GKP qubits. Alternatively, any lattice can be constructed with passive interferometers starting with GKP Bell pairs, where the number of physical modes per macronode is the valency of the underlying canonical graph state and the GKP Bell pairs are distributed to match the entanglement in the canonical graph state. Relating measurements on the physical modes with effective measurements on the underlying lattice requires a list of corrections known as a dictionary protocol Alexander et al. (2014), which we describe in Appendix XI.6. Note, that in the following we use the terms GKP state and sensor state interchangeably, always meaning either a sensor state or one half of a GKP Bell pair formed by interfering sensor states.
We wish to design a modular construction of the required lattice. To this end, we identify a unit cell consisting of 24 sensor states or 12 GKP Bell pairs that can be tiled to construct a macronode RHG lattice of arbitrary size. From this unit cell, four GKP nodes are sent to neighboring unit cells, and correspondingly four are received. Two of the modes are delayed to form temporal links. This is illustrated in Fig. 11. As the primal and dual layers are created in a single time-step, we avoid the requirement for adaptive optics that have often been used to account for the different entanglement structure in primal and dual layers Bourassa et al. (2021); Larsen et al. (2021). To teleport non-Clifford operations onto logical qubits within the surface code, physical magic states are injected into the RHG lattice Litinski (2019a); Herr et al. (2018). The physical magic state is interfered with a vacuum state at a beamsplitter and is used in place of one GKP Bell pair in the RHG macronode.
VIII FTQC Threshold Simulation
Several CV FTQC architectures based on GKP qubits and the surface code exist in the literature and have been characterised by their GKP effective squeezing threshold. Refs. Fukui et al. (2018) and Bourassa et al. (2021) used constructions based on $C_{Z}$ gates and found thresholds of $10$ dB and $10.5$ dB respectively, however these ignored the additional noise from finite squeezing in physical gates. In Ref. Larsen et al. (2021) found a threshold of $12.7$ dB while accounting for finite squeezing. Ref. Tzitrin et al. (2021) proposed a macronode-based architecture which did away with the need for $C_{Z}$ gates and found a threshold of $10.1$ dB. Refs. Bourassa et al. (2021); Larsen et al. (2021); Tzitrin et al. (2021) all require fast reconfigurable optics either in the computation or in the preparation of GKP states with a sufficiently high success probability.
In the following section, we analyze the fault-tolerant performance of our architecture. In order to allow comparison with other architectures, we first use a phenomenological error model used widely in the literature Fukui et al. (2018, 2021); Larsen et al. (2021), which models realistic GKP states as ideal GKP states acted on by a Gaussian displacement channel. Formally, this is represented as Tzitrin et al. (2021)
$\mathcal{E}\left(\rho,\Delta\right)=\frac{1}{\sqrt{\pi}\Delta^{2}}\int\mathrm{d}\mathbf{\alpha}\exp\left(-\frac{1}{\Delta^{2}}\left|\mathbf{\alpha}\right|^{2}\right)\hat{D}\left(\mathbf{\alpha}\right)\rho\hat{D}^{\dagger}\left(\mathbf{\alpha}\right)$ (1)
Which has the effect of replacing the delta function teeth of an ideal GKP state with Gaussian distributions with a standard deviation, $\Delta$,where $\Delta$ describes the same broadening of the GKP teeth as $\Delta_{D}$ described in Sec. II Bourassa et al. (2021). We can link this to the GKP effective squeezing in dB as $S_{\mathrm{dB}}=-\log_{10}\left(\Delta^{2}\right)$ Bourassa et al. (2021). Despite its simplicity, this error model can describe many common noise sources such as finite squeezing and photon loss Tzitrin et al. (2021); Bourassa et al. (2021); Larsen et al. (2021); Noh and Chamberland (2020). We then move on to quantify the fault-tolerance performance in terms of the input Gaussian resource for PhANTM - the squeezing level of the multi-mode cluster state. The input for these simulations are distributions of GKP squeezings obtained from simulations of the proposed GKP generation pipeline.
In both cases, to determine the logical error rate we track how the GKP effective squeezing propagates through the optical circuit used to generate the macronode lattice and use the propagated noise variances to simulate noisy homodyne measurements. We apply the corresponding post-processing to reduce the physical macronode measurements to the effective noisy measurements on the canonical lattice, as shown in Appendix XI.6. With these noisy measurements we determine the probability of a qubit level error on each node of the canonical lattice, see Appendix XI.7 for more detail. These probabilities are used to sample a set of errors and generate the corresponding syndrome of the RHG lattice. This syndrome is then decoded using the pyMatching implementation of the minimum weight perfect matching decoder Higgott and Gidney (2023). This simulation method is described in more detail in Appendix XI.8 and is similar to the method used in Ref Bourassa et al. (2021), adapted to the macronode approach in Ref. Tzitrin et al. (2021) and for our specific GKP generation method.
VIII.1 Phenomenological error model
Here, we present results which tailor the phenomenological noise model to the PhANTM and adaptive breeding approach to generate GKP states. As described in Sec. V, generating GKP states from PHANTM and breeding results in GKP states that are inherently asymmetric due to the different physical processes that determine the squeezing in the GKP quadratures. Note that this is equivalent to grid states with asymmetric spacing in the two quadratures Stafford and Menicucci (2023) which has also been studied in Walshe et al. (2024). This feature was incorporated into our simulation method by defining $p$ and $q$ noise variances separately for each GKP state. While we can also set each GKP state independently, for simplicity we assume that that the $p$ and $q$ noise are identical for each GKP state. To determine the 2D threshold curve for unbalanced quadrature noise we employ the method used in Ref. Bartolucci et al. (2021). By expressing the $p$ and $q$ noise in terms of a dummy variable, $\Delta_{p}=c_{p}\times x$, $\Delta_{q}=c_{q}\times x$, we fix the ratio of $p$ and $q$ noise via $c_{p}/c_{q}$ and use $x$ to find the threshold for a given ratio.
Figure 12: (a) 2D threshold plot for independent noise in the GKP quadratures, in terms of decibel squeezing, $S_{\mathrm{dB}}=-10\log\left(\Delta^{2}\right)$. Combinations of squeezing values above the line are correctable and those below are not. Dotted lines show the marginals with infinite squeezing in one of the quadratures. The inset shows an example threshold plot for a single point, shown in blue, in the main panel. Each point is averaged $10^{5}$ times. (b) Plot showing the dependence of the squeezing threshold on the distribution of GKP effective squeezing. On each Monte Carlo run, first the GKP effective squeezing for each mode from a normal distribution with standard deviation $\sigma_{d}$ before the noisy homodyne values are generated from the sampled squeezing value.
Fig. 12 shows the resultant surface, where we have shaded the correctable region. The surface is symmetric with marginal squeezing thresholds of $7.2\mathrm{dB}$. With balanced squeezing in each quadrature we recover a threshold of $10.2\mathrm{dB}$, which is inline with similar error models Tzitrin et al. (2021).
A second adaption stems from the fact that the size of the cat states from PhANTM depend on the number of photons subtracted and so are inherently probabilistic. This results in GKP states that do not have a single squeezing value for each quadrature, but rather are described by a distribution of effective squeezing. To include this in our simulation framework we employ a two-stage sampling procedure where for each mode and for each Monte Carlo iteration we sample first the squeezing variance for each mode from a Gaussian distribution with standard deviation given by $\sigma_{d}$ and then we sample the noisy homodyne value. Fig. 12(b) shows how the squeezing threshold depends on $\sigma_{d}$.
VIII.2 End-to-end simulations
As in many similar architectures Larsen et al. (2021); Fukui et al. (2018); Chamberland and Cross (2019); Bourassa et al. (2021); Tzitrin et al. (2021), the previous section abstracted the thresholds in terms of GKP effective squeezing. While this allows a comparison against different architectures, it does not give an estimate of the physical requirements for our approach. A key resource for our architecture is the squeezing level of the multi-mode Gaussian cluster state. In this section we aim to determine the threshold in terms of this resource. To this end we perform simulations of PhANTM and adaptive breeding and then use the resultant GKP states as the inputs to the logical-level simulations. We fix the PhANTM and breeding setup to use 8 PNR detectors per frequency mode with gradient beamsplitters, see Sec. IV. We chose three breeding rounds and employ a threshold for cat squeezing of $r_{lb}=0.5$ ($4.34\mathrm{dB})$, below which a momentum squeezed state from the cluster state is swapped in, as described in Sec. V. We then sweep the cluster squeezing and simulate the logical error rate for several code distances, repeating this simulation for a range of PhANTM steps.
Figure 13: Fault-tolerant threshold for cluster squeezing for a range of PhANTM step numbers. Bottom panels show the corresponding mean and standard deviations of GKP effective squeezing distributions input to the logical-level simulation. These results used $10^{3}$ PhANTM iterations, $10^{3}$ adaptive breeding iterations and $10^{5}$ logical error rate trials.
We find cluster squeezing thresholds of $12.1,12.7,13$ dB for $20$, $15$ and $10$ PhANTM steps. We see that the required cluster squeezing increases as the number of PhANTM steps, and therefore the mean cat state amplitude $\alpha$, decreases. The lower panels in Fig. 13 show the mean and standard deviation of the GKP effective squeezing distributions output from the breeding process. We also compare the adaptive breeding protocol, used in Fig. 13, to standard breeding, where cat states must be squeezed exactly to the GKP spacing. With standard breeding, no threshold was found below $13.5$ dB, indicating that adaptive breeding can be used to reduce the threshold by at least $1.5$ dB and possibly more. While our simulation takes into account all generated GKP states, not just those above threshold, we can quantify the proportion of GKP states that fall above threshold, allowing us to not only compare the Gaussian squeezing resources required but also the generation probability with other GKP generation protocols. To this end, we determine the likelihood that a GKP state drawn randomly from the generated distribution has a squeezing level above the threshold, that is the combination of squeezing in each quadrature puts it in the ’correctable’ region of Fig. 12. This likelihood can then be taken as the success probability of our protocol. Some examples are shown in table 1.
Cluster squeezing (dB) PhANTM steps
10 20
12.0 0.26 0.45
13.0 0.56 0.91
Table 1: Success rates for which adaptive breeding produces sensor states with squeezing levels above the FT threshold for different levels of cluster squeezing and PhANTM steps.
While direct simulations of the required Gaussian squeezing thresholds for other GKP generation methods are missing from the literature, we can compare to several approaches that benchmark themselves by generating $10$ dB GKP states, which are at the threshold for several fault-tolerant architectures Tzitrin et al. (2021); Fukui (2023); Fukui et al. (2018). The highest success probability for single-shot GKP generation protocols are found in Ref. Takase et al. (2024); however they require input $18-20$ dB of Gaussian squeezing, and photon number resolution up to 40 photons. On the other hand, lower input squeezing resources can be used as in Ref. Tzitrin et al. (2020), requiring $12$ dB, and Ref. Winnel et al. (2024) is as low as $6\mathrm{dB}$. However, in the former, success probabilities are on the order of $10^{-5}$ and the latter requires the generation of large, $n=10$, Fock states which are experimentally challenging .
IX Discussion
We have presented an architecture for CV FTQC that uses the time-frequency-space degrees of freedom to create GKP states above the FT threshold with high probabilities, creates distillable magic states with a supply of realistically squeezed GKP states and cat states, and can be implemented on low-loss photonic chips in a modular architecture. This work proposes a static state generation process that produces a distribution of GKP effective squeezing, affecting the FT threshold only modestly compared to idealized GKP states with a set squeezing target, as shown in Fig. 12. The results presented in the paper require Gaussian squeezing levels of $12$ dB to $13$ dB. Low photon number resolution is needed for cat state generation after which all steps require homodyne detections and linear optical CV operations. The architecture can be implemented without on-chip switches or quantum memories, resulting in a fully-passive architecture.
The experimental hardware needed for the architecture is ambitious but feasible. The squeezing level needed is below the state-of-the-art levels achieved in free-space Vahlbruch et al. (2016). In recent years, the squeezing measured from on-chip sources has steadily improved, reaching $8.3$ dB Kashiwazaki et al. (2023), with efforts underway to reach levels seen in free-space. The other major resource needed in our approach is photon number resolving detection. While TES devices offer resolutions in excess of $100$ photons Eaton et al. (2022b) and have traditionally been used in CVQC, SMSPDs Kong et al. (2024) with low data acquisition requirements operating at higher temperatures have been shown to detect up to $10$ photons, sufficient for our architecture. Simultaneously, chip-integrated room-temperature PNR technologies capable of detecting up to $10$ photons at MHz scales are also in development Nehra et al. (2019). Finally, while losses in photonic circuit components need to be minimized for implementing our architecture on photonic chips, Silicon nitride platforms currently provide some of the lowest losses available for passive components Team (2024) at levels that look promising for FTQC.
Multiple improvements, optimizations, and validations of the presented architecture remain to be explored. Higher probability methods for cat state generation via optimizing the squeezing and photon number resolution are possible. Adaptive breeding, used in both sensor state and magic state generation protocols, can be optimized with advanced adaptation algorithms. In the case of magic states, targeting a higher success probability while simultaneously reducing the quality of GKP states needed for magic generation looks possible with further development of the CV-QEC process with non-Gaussian states. The squeezing requirements of the architecture could also be lowered by moving to cluster states with a lower valency Paesani and Brown (2023) and employing more sophisticated GKP decoders, such as the ones taking into account correlations introduced by linear optics Walshe et al. (2024). While the effect of detection losses in PhANTM has been investigated Eaton et al. (2022a), estimating the loss and noise thresholds to target for hardware development would require more accurate and comprehensive error models, including the effects of Gaussian noise on effective squeezing and optical loss on Wigner negativity at each step of our architecture.
While many improvements could be made and the performance under losses are yet to be fully quantified, we believe that the modular and passive architecture for photonic CV FTQC presented in the paper will constitute a foundational design for photonic quantum computers above the FT threshold in the future.
X Acknowledgements
Part of the QEC simulation code was funded under DARPA Contract HR00112390051. The views, opinions, and/or findings expressed are those of the author(s) and should not be interpreted as representing the official views or policies of the Department of Defense or the U.S. Government.
We acknowledge insightful conversations with QC82 co-founders Olivier Pfister, Xu Yi, Joe Campbell, and Andreas Beling on theoretical and experimental CVQC, squeezing sources, homodyne detection, photon number resolution, and many other topics. We are thankful to Leandre Brunel for sharing some of his work on GKP lattice construction methods.